Date:
Wednesday, January 28, 2026
Telescope:
CTIO Blanco 4-m
Submitter:
Sean Points
Submitter Email:
Observer Support:
Manuel Hernandez
Night Assistant:
Claudio Aguilera
Also Present:
R Hinojosa
Problems:
When we stated the guider tests near the horizon, we went to a bright star to redo the telescope pointing. After that, we performed a focus sequence. There was a problem with M2 and the focus value did not change. We then noticed that the focus value, which was initially ~18950 units was decreasing on the GUI. It reached a value of ~300.
Manuel H was able to re-initialize M2 and we moved to zenith and restored the focus value to 18950. This occurred between 23:00-24:00 local time. Can we please check the logs to see what happened.
Program:
Proposal:
Newfirm Engineering
Target of Oportunity:
No
Time Observed:
0.0
Time Engineering:
8.5
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Description:
NEWFIRM Engineering
Primary Investigator:
Points
Email:
Institution:
CTIO
Instrument:
CTIO 4m Newfirm
Organization:
NOAO
Total Program Time:
8.5
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2026-01-28 | 0.0 | 8.5 | 0.0 | 0.0 | 0.0 | 8.5 |
Guider tests (no guiding for 180s and guiding for 180s) at zenith. No problems.
Guider tests (no guiding for 300s and guiding for 300s) near horizon. No problems.
Donuts of bright stars at Dec = [-30, -15, +05, +15 deg]. No problems.
Focus offset tests wth filters. No problems.